# First Hitting Times on a Quantum Computer: Topological Effects and Dark States

### 24 Feb 2024 v1

@@ IV.1 On-site protocol (experiment)
Fig. 3 (a) shows the classical simulation of the mean return time as a function of $\gamma\tau$ and $\alpha$ for $N=20$. In comparison to Fig. 2 the areas for $\langle n\rangle=1$ (dark blue) and $\langle n\rangle=2$ (medium blue) are broadened. For the experimental computation of $\langle n\rangle$ we utilize the quantum circuit in Fig. 3 (b). The on-site protocol can be emulated by measuring the upper qubit after each unitary $U$.
@@
Let us first discuss the results of the asymptotic theory using Eq. (7). In Fig. 3 (c) - (f), we consider $\langle n\rangle$ versus the dimensionless sampling time $\gamma\tau$, as well as the magnetic flux $\alpha$ (black solid line). For almost any $\tau$, on average, we return after two measurements, i.e., $\langle n\rangle=2$. This holds when the magnetic flux is turned off ($\alpha=0$). In contrast, when $\alpha\neq 0$ we have besides very special choices of $\tau$, $\langle n\rangle=3$ which is related to the fact that the magnetic flux $\alpha>0$ removes the degeneracy by breaking the time-reversal symmetry and the rotational invariance. In addition, one observes typical sudden plunges in $\langle n\rangle$. These are found for special values of $\gamma\tau$ and $\alpha$. As mentioned in the Introduction, these transitions, e.g., dips in $\langle n\rangle$ when plotted versus $\gamma\tau$ or $\alpha$ are related to a discontinuous change of the topology of the underlying generating function [5, 10].
@@
Quantization of the mean return time of the experimental result (light blue circles) is clearly evident for most values of $\gamma\tau$ and $\alpha$. Far from transitions in $\langle n\rangle$, theory predicts, and the experiments confirm, that for on-site measurements, with and without magnetic flux, $\langle n\rangle\approx 2$ [Fig. 3 (c)] and $\langle n\rangle\approx 3$ [Fig. 3 (d)-(f)], respectively. Notwithstanding these successes, theory ($N\rightarrow\infty$) and experiment depart close to the parameters in which the broadening of the dark blue and medium blue areas is obvious in Fig. 3 (a). When we choose path (III), in the ($\alpha$, $\gamma\tau$) plain, we are in the vicinity of a phase factor matching curve, though strictly not exactly on it. Since the number of measurements $N$ is finite in the experimental study, we are effectively witnessing a topological state with $\langle n\rangle=2$, instead of $\langle n\rangle=3$ predicted by theory, see Fig. 3 (e). We observe not only a broadening of the resonances, but also additional resonances for finite $N$ that are not present in the theoretical graph ($N\rightarrow\infty$) [Fig. 3 (d)]. Finite resolution comes with a price. This is relevant when triplets of dips of $\langle n\rangle$ are in the vicinity of one another on path (II). Indeed, when $\alpha=0.5$, we see from the theoretical prediction in Fig. 3 (d) three nearby dips (black solid line), while in reality, we see one wide resonance (light blue circles). Thus, the finite resolution wipes out the fine details of the model because all three phase factors are almost matching.
@@
The cause of the deviation between theory (black solid lines) and experiment (light blue circles) close to these topological transitions is not entirely clear. It might be due to noise, readout errors or finite resolution effects. Without a doubt, all of these effects may play some role. However, for the mean return time under study, we concluded that the main factor is the finite resolution of the experiment. The unitary dynamics and measurements implemented on the quantum computer work well on the time and size scale of our experiment, indicated by the fact that the experimental points lie close to the simulation on a classical computer (blue dashed lines). Finite resolution effects are discussed in detail in Sec. VI.
Figure 3: (a) Simulation of the mean return time $\langle n\rangle$ for $N=20$ for the on-site protocol as a function of $\gamma\tau$ and $\alpha$. For $\langle n\rangle=1$ three phase factors match (dark blue), for $\langle n\rangle=2$ two phase factors match (medium blue) and for $\langle n\rangle=3$ all phase factors are different (white). The paths through the parameter space labeled as (I), (II), (III) and (IV) are indicated with red dashed lines and correspond to (c), (d), (e) and (f), respectively. Compared to the phase factor matching diagram in Fig. 2 the lines are broadened due to finite $N$. (b) Quantum circuit for two qubits representing the three localized states for the on-site protocol with the initial state |$\ket{\bf{0}}=\ket{01}$, where the unitary $U=\exp{(-iH\tau)}$ and measurements are applied alternately. Only the upper qubit is measured, since we need to detect the state |$\ket{\bf{0}}$ while we do not want to receive the information that allows us to distinguish the states |$\ket{\bf{1}}$ and |$\ket{\bf{2}}$. (c) - (f) Mean return time $\langle n\rangle$ recorded on IBM Sherbrooke (light blue circles), obtained with the asymptotic theory for $N\rightarrow\infty$ (black solid lines) and simulated for $N=20$ (blue dashed lines): (c) [path (I)] $\langle n\rangle$ versus $\gamma\tau$ for $\alpha=0$. (d) [path (II)] When $\alpha=0.5$ we almost always find $\langle n\rangle=3$. As explained in the text, this is related to the removal of energy-level degeneracy when the magnetic flux is turned on. Note that some fine structure details predicted by the asymptotic theory, for $\alpha=0.5$, are washed out in the experiment, see the three nearby dips that merge into one resonance here. (e) [path (III)] Mean return time $\langle n\rangle$ as a function of $\alpha$ for $\gamma\tau=3\pi/\sqrt{3}$. $\langle n\rangle$ develops a plateau, as well as additional transitions that are absent for $N\rightarrow\infty$. (f) [path (IV)] Mean return time $\langle n\rangle$ for $\gamma\tau=2\pi/\sqrt{3}$ as a function of $\alpha$.
Figure 4: (a) Simulation of the mean return time $\langle n\rangle$ for $N=20$ for the tracking protocol as a function of $\gamma\tau$ and $\alpha$. For $\langle n\rangle=1$ three phase factors match (dark blue), and for $\langle n\rangle=3$ all phase factors are different (white). Paths through the parameter space labeled (I), (II), (III) and (IV) are indicated with red dashed lines and correspond to (c), (d), (e), and (f), respectively. Compared to the phase factor matching diagram in Fig. 2 the dark blue areas are broadened due to finite $N$. (b) Quantum circuit for two qubits representing the three localized states for the tracking protocol with the initial state |$\ket{\bf{0}}=\ket{00}$, where the unitary $U=\exp{(-iH\tau)}$ and the measurements are applied alternately. (c) - (f) Mean return time $\langle n\rangle$ recorded on IBM Sherbrooke (light blue circles), obtained with the asymptotic theory for $N\rightarrow\infty$ (black solid line) and simulated for $N=20$ (blue dashed line): (c) [path (I)] $\langle n\rangle$ versus $\gamma\tau$ for $\alpha=0$. $\langle n\rangle$ is quantized ($\langle n\rangle=3$) and its value, for almost any choice of $\gamma\tau$, differs by unity from the on-site measurement protocol, the case presented in Fig. 3 (c) when $\langle n\rangle=2$. (d) [path (II)] For finite magnetic flux $\alpha=0.5$ the transitions to $\langle n\rangle=1$ are absent except in the Zeno limit at $\gamma\tau=0$ for $N\rightarrow\infty$. This is different for $N=20$. The result for $\alpha=0.5$ shows a transition to approximately $\langle n\rangle=1$ at $\gamma\tau=3.63$. (e) [path (III)] Mean return time $\langle n\rangle$ for $\gamma\tau=3\pi/\sqrt{3}$. For $N=20$ $\langle n\rangle$ develops additional transitions that are absent for $N\rightarrow\infty$ (f) [path (IV)] Mean return time $\langle n\rangle$ for $\gamma\tau=2\pi/\sqrt{3}$ as a function of $\alpha$.
@@ IV.2 Tracking protocol (experiment)
Fig. 4 (a) shows the classical simulation of the mean return time as a function of $\gamma\tau$ and $\alpha$ for $N=20$. In comparison to Fig. 2 the areas for $\langle n\rangle=1$ (dark blue) are broadened and the lines where two phase factors match are irrelevant. For the experimental computation of $\langle n\rangle$ we implement the quantum circuit in Fig. 4 (b), where both qubits are measured after applying the unitary $U$.
We start by discussing the theory for $N\rightarrow\infty$ (black solid line). The tracking protocol gives with and without flux $\langle n\rangle=3$, according to theoretical predictions. Note that any small deviation of $\tau$ from the special revival times or $\tau=0$, that is, the Zeno limit, yields $\langle n\rangle=3$ as shown in Fig. 4 (c) - (f) as a black solid line. For the finite magnetic flux $\alpha=0.5$ in Fig. 4 (d), there are no special sampling times $\tau$ except $\tau=0$. (Similar for $\gamma\tau=3\pi/\sqrt{3}$ and finite $\alpha$ in Fig. 4 (e)). In the vicinity of the Zeno limit, when revivals play a role, or more generally, when three phase factors merge, the deviations between the theory valid for $N\to\infty$ and experiments (light blue circles) for $N=20$ are large. We obtain a large deviation for the parameters in which the broadening of the dark blue areas is observed in Fig. 4 (a). The broadening is present in direction of $\gamma\tau$ and $\alpha$ which leads not only to a broadening of the resonances, but also to additional resonances for finite $N$ that are not present in the theoretical graph ($N\rightarrow\infty$) [Fig. 4 (d) and (e)] but in the finite $N$ simulation (blue dashed line). Similarly to the on-site protocol, the deviation of the experimental values from the asymptotic theory is due to finite-resolution effects. These are discussed in detail in Sec. VI.
@@ V Dark States on IBM quantum computers
Figure 5: Detection probability $0\leq P_{{\rm det}}\leq 1$ as a function of the sampling time $\gamma\tau$ and the phase $\phi$ as well as $\gamma\tau$ and magnetic flux $\alpha$ obtained from IBM Sherbrooke with $N=10$ for the on-site and tracking protocol. The dark blue color corresponds to $P_{\rm det}=0$ (dark state), while the white color corresponds to $P_{\rm det}=1$ (bright state). (a) $P_{{\rm det}}$ as a function of $\gamma\tau$ and $\phi$ for the on-site protocol. Dark states are found when $\phi=\pi$ is the result of destructive interference and for $\gamma\tau\to 0$ due to the Zeno effect. Additional horizontal bands show dark states, which are found when the time of revival in the initial state is the same as the sampling time $\gamma\tau$. (b) $P_{\rm det}$ as a function of $\gamma\tau$ and $\phi$ for the tracking protocol. The dark states found for $\phi=\pi$ for the on-site protocol turn bright when the tracking method is used, since the latter breaks spatial symmetry, while the former does not. The horizontal dark bands are due to revivals that forbid detection of the system in the target state. (c) and (d) $P_{\rm det}$ as a function of $\gamma\tau$ and magnetic flux $\alpha$. The on-site protocol (c) and the tracking protocol (d) exhibit vastly different trends, i.e., dark states are present in the tracking protocol where three phase factors match and in the on-site protocol where two and three phase factors match. As explained in the text, (c) yields the phase factor matching diagram shown in Fig. 2, while (d) corresponds to three phase factors matching, which are indicated by blue circles in Fig. 2.
@@ V.2 Dark states for a finite magnetic flux
Finally, we pose the question of whether the total detection probability changes if we choose |$\ket{\psi}_{0}=\ket{\bf{2}}$. The total detection probability for the transition from |$\ket{\bf{1}}$ to |$\ket{\bf{0}}$ and |$\ket{\bf{2}}$ to |$\ket{\bf{0}}$, respectively, is different and depends on the direction of the magnetic field for $N=10$. The difference of the total detection probabilities is shown in Fig. 6, where an asymmetry is observed around the broadened phase-matching areas and lines. The total detection time of the chiral quantum walker for the transfer depends on the direction (clockwise or counter clockwise) and also on the sign of the magnetic flux. Breaking of the time-reversal symmetry can therefore enhance or suppress the total detection and therefore provides a directional control for certain parameters.
@@ VI Finite resolution
By finite resolution, we mean that measurements are performed $N$ times per run. Typically, we record the target state for the first time in some random measurement attempt numbered $n_{i}\leq N$. We repeat this process by constructing an ensemble of trajectories to sample the mean return time which is maximally the number of runs on the quantum computer that we choose to be $32,000$. We have already explained that in some cases we might not find the target within these $N$ measurements. Hence, in the figures presented, we show sample averages that, in principle, depend on our choice of $N$. More specifically, we define $K\leq 32,000$ as the number of paths for which we record the target state. Then, the sample mean $n$ conditioned on the target measurement state is $\langle n\rangle=\sum_{i=1}^{K}n_{i}\big{/}K$ where we exclude the cases that we did not detect. Far from the topological transitions, the option of zero detection per run (null measurement) of size $N$, is rare, and then the conditional measurement and the theoretical predictions yield similar results. However, as mentioned, close to topological transitions, we see deviations, which are due to the finite value of $N$. It should be noted that these deviations are a blessing in the sense that a precise measurement of a jump of $\langle n\rangle$, for a special value of $\tau$, as presented as black solid lines in Figs. 3 and 4, is non-physical, as the width of the transition cannot be zero. In the asymptotic theory ($N\rightarrow\infty$) for a special value of model parameters (here $\gamma\tau$ and $\alpha$), $\langle n\rangle$ exhibits a pointwise discontinuity related to a jump in the winding number and the dark lines in Fig. 5 become infinitely thin.
@@ VI.1 Broadening effect
We present an exact solution to the problem related to our experimental setup for both on-site and tracking measurement protocols. Previously, the broadening effect was studied in more generality for the on-site measurements, although the asymptotic behavior of large $N$ was studied there [35].
VI.2 On-site protocol (broadening)
We first consider the broadening effect using the on-site measurement protocol and $\alpha=0$. Let $F_{n}$ be the probability of finding the target for the first time at the $n-$th measurement. The case under study is exactly solvable [10, 36]. Specifically, with the eigenvalues of Eq. (3) we obtain
@@ VI.2 On-site protocol (broadening)
where $|z|^{2}=5/9+4/9\cos(3\gamma\tau)$. Here, we stick to the notation in [10], where in general, $z$ stands for the zeros of the generating function. When $|z|^{2}=1$, we have $F_{1}=1$ and then $\langle n\rangle=1$. Otherwise, $\langle n\rangle=\sum_{n=1}^{\infty}nF_{n}=2$, which is a behavior that was presented in Fig. 3 (c). Clearly, when $|z|^{2}$ is slightly less than one, so is $F_{1}$, yet from the recurrence theorem we have $\lim_{N\rightarrow\infty}\langle n\rangle(N)=2$ implying a very slow decay of $F_{n}\propto|z|^{2(n-2)}$ for $n>1$. Physically, close to $|z|^{2}=1$ we usually detect the system with the first click, but in rare events, we click no. Then the wave function does not overlap with the detected state, and recording a yes becomes unlikely. On average we find that the return click yes after $\langle n\rangle=2$ attempts.
@@
We will study this expression in the vicinity of the revival times, namely when $|z|^{2}\to 1$, where we also include the Zeno limit in this category. Note that if $|z|^{2}=1$ we have $\langle n\rangle(N)=1$ for any $N$.
We note that the expressions in Eqs. (11) and (12) have the same form for a two-level system, except for a different function of $|z|^{2}$ [36]. This indicates a universal behavior of $F_{n}$ for finite $N$ and their related expectation values.
@@
Recall that close to the transition point $\langle n\rangle=1$ for any large though finite $N$. On the other hand, if we choose a control parameter like $\gamma\tau$ to be far from the transition, then $\langle n\rangle\approx 2$, since $N$ is large.
VI.3 Tracking protocol (broadening)
Now let us consider the broadening effect using the tracking protocol and $\alpha=0$. The probability of finding the quantum walker the first time is for the tracking protocol

### 10 Apr 2024 v2

@@ IV.1 On-site protocol (experiment)
Fig. 3 (a) shows the parameter regime (red dashed line) utilized for the computation of the mean return time as well as the classical simulation of the mean return time as a function of $\gamma\tau$ and $\alpha$ for $N=20$. In comparison to Fig. 2 the areas for $\langle n\rangle=1$ (dark blue) and $\langle n\rangle=2$ (medium blue) are broadened. The figure clearly shows how the conditions for phase factor matching can lead to jumps in $\langle n\rangle$.
For the experimental computation of $\langle n\rangle$ we utilize the quantum circuit in Fig. 3 (b). The on-site protocol can be emulated by measuring the upper qubit after each unitary $U$.
@@
Let us first discuss the results of the asymptotic theory using Eq. (7). In Fig. 3 (c) - (f), we consider $\langle n\rangle$ versus the dimensionless sampling time $\gamma\tau$, as well as the magnetic flux $\alpha$ (black solid line). For almost any $\tau$, on average, we return after two measurements, i.e., $\langle n\rangle=2$. This holds when the magnetic flux is turned off ($\alpha=0$). In contrast, when $\alpha\neq 0$ we have besides very special choices of $\tau$, $\langle n\rangle=3$ which is related to the fact that the magnetic flux $\alpha>0$ removes the degeneracy by breaking the time-reversal symmetry and the rotational invariance. In addition, one observes typical sudden plunges in $\langle n\rangle$. These are found for special values of $\gamma\tau$ and $\alpha$. As mentioned in the Introduction, these transitions, e.g., dips in $\langle n\rangle$ when plotted versus $\gamma\tau$ or $\alpha$ are related to a discontinuous change of the topology of the underlying generating function [5, 6, 11].
@@
Quantization of the mean return time of the experimental result (light blue circles) is clearly evident for most values of $\gamma\tau$ and $\alpha$. Far from transitions in $\langle n\rangle$, theory predicts, and the experiments confirm, that for on-site measurements, with and without magnetic flux, $\langle n\rangle\approx 2$ [Fig. 3 (c)] and $\langle n\rangle\approx 3$ [Fig. 3 (d)-(f)], respectively. Notwithstanding these successes, theory ($N\rightarrow\infty$) and experiment depart close to the parameters in which the broadening of the dark blue and medium blue areas is obvious in Fig. 3 (a). When we choose path (III), in the ($\alpha$, $\gamma\tau$) plain, we are in the vicinity of a phase factor matching curve, though strictly not exactly on it. Since the number of measurements $N$ is finite in the experimental study, we are effectively witnessing a topological state with $\langle n\rangle=2$, instead of $\langle n\rangle=3$ predicted by theory, see Fig. 3 (e). We observe not only a broadening of the resonances, but also additional resonances for finite $N$ that are not present in the theoretical graph ($N\rightarrow\infty$) [Fig. 3 (d)].
Furthermore, experiment and asymptotic theory do not match when triplets of dips of $\langle n\rangle$ are in the vicinity of one another on path (II). Indeed, when $\alpha=0.5$, we see from the theoretical prediction in Fig. 3 (d) three nearby dips (black solid line), while in reality, we see one wide resonance (light blue circles). Thus, fine details are wiped out near the point where all three phase factors are almost matching.
@@
What is the cause of the deviation between theory (black solid lines) and experiment (light blue circles) close to these topological transitions? It might be due to noise, readout errors, or finite resolution effects. Without a doubt, all of these effects may play some role. However, for the mean return time under study, we concluded that the main factor is the finite resolution of the experiment. The unitary dynamics and measurements implemented on the quantum computer work well on the time and size scale of our experiment, indicated by the fact that the experimental points lie close to the simulation on a classical computer (blue dashed lines). Finite resolution effects are discussed in detail in Sec. VI.
@@ IV.2 Tracking protocol (experiment)
Fig. 4 (a) shows the parameter regime (red dashed line) utilized for the computation of the mean return time as well as the simulation of the mean return time as a function of $\gamma\tau$ and $\alpha$ for $N=20$ for the tracking protocol. Compared to Fig. 2 the areas for $\langle n\rangle=1$ (dark blue) are broadened and in comparison to the on site protocol Fig. 3 (a) we observe the dark color ($\langle n\rangle$) more sporadically, i.e., the lines where only two phase factors match are irrelevant. For the experimental computation of $\langle n\rangle$ we implement the quantum circuit in Fig. 4 (b), where both qubits are measured after applying the unitary $U$.
We start by discussing the theory for $N\rightarrow\infty$ (black solid line). The tracking protocol gives with and without flux ($\alpha$) $\langle n\rangle=3$, for nearly all $\tau$, according to theoretical predictions. Note that any small deviation of $\tau$ from the special revival times or $\tau=0$, that is, the Zeno limit, yields $\langle n\rangle=3$ as shown in Fig. 4 (c) - (f) as a black solid line. For the finite magnetic flux $\alpha=0.5$ in Fig. 4 (d), there are no special sampling times $\tau$ except $\tau=0$. (Similar for $\gamma\tau=3\pi/\sqrt{3}$ and finite $\alpha$ in Fig. 4 (e)). When three phase factors merge, the deviations between the theory valid for $N\to\infty$ and experiments (light blue circles) for $N=20$ are large. We obtain a large deviation for the parameters in which the broadening of the dark blue areas is observed in Fig. 4 (a) between experiment and assymptotic theory. The broadening is present in direction of $\gamma\tau$ and $\alpha$ which leads not only to a broadening of the resonances, but also to additional resonances for finite $N$ that are not present in the theoretical graph ($N\rightarrow\infty$) [Fig. 4 (d) and (e)] but in the finite $N$ simulation (blue dashed line). Similarly to the on-site protocol, the deviation of the experimental values from the asymptotic theory is due to finite-resolution effects. These are discussed in detail in Sec. VI.
@@ V.1 Dark states for zero magnetic flux
Figure 5: Detection probability $0\leq P_{{\rm det}}\leq 1$ as a function of the sampling time $\gamma\tau$ and the phase $\phi$ as well as $\gamma\tau$ and magnetic flux $\alpha$ obtained from IBM Sherbrooke with $N=10$ for the on-site and tracking protocol. The dark blue color corresponds to $P_{\rm det}=0$ (dark state), while the white color corresponds to $P_{\rm det}=1$ (bright state). (a) $P_{{\rm det}}$ as a function of $\gamma\tau$ and $\phi$ for the on-site protocol. Dark states are found when $\phi=\pi$ is the result of destructive interference and for $\gamma\tau\to 0$ due to the Zeno effect. Additional horizontal bands show dark states, which are found when the time of revival in the initial state is the same as the sampling time $\gamma\tau$ and $\alpha=0$. (b) $P_{\rm det}$ as a function of $\gamma\tau$ and $\phi$ ($\alpha=0$) for the tracking protocol. The dark states found for $\phi=\pi$ for the on-site protocol turn bright when the tracking method is used, since the latter breaks spatial symmetry, while the former does not. The horizontal dark bands are due to revivals that forbid detection of the system in the target state. (c) and (d) $P_{\rm det}$ as a function of $\gamma\tau$ and magnetic flux $\alpha$ ($\phi=0$). The on-site protocol (c) and the tracking protocol (d) exhibit vastly different trends, i.e., dark states are present in the tracking protocol where three phase factors match and in the on-site protocol where two and three phase factors match. As explained in the text, (c) yields the phase factor matching diagram shown in Fig. 2, while (d) corresponds to three phase factors matching, which are indicated by blue circles in Fig. 2.
@@ V.2 Dark states for a finite magnetic flux
Finally, we pose the question of whether the total detection probability changes if we choose |$\ket{\psi}_{0}=\ket{\bf{2}}$. The total detection probability for the transition from |$\ket{\bf{1}}$ to |$\ket{\bf{0}}$ and |$\ket{\bf{2}}$ to |$\ket{\bf{0}}$, respectively, is different and depends on the direction of the magnetic flux $\alpha$ for $N=10$. The difference of the total detection probabilities is shown in Fig. 6, where an asymmetry is observed around the broadened phase-matching areas and lines. The total detection time of the chiral quantum walker for the transfer depends on the direction (clockwise or counter clockwise) and also on the sign of the magnetic flux. Breaking of the time-reversal symmetry can therefore enhance or suppress the total detection and therefore provides a directional control for certain parameters.
@@ VI Finite resolution
Finite resolution plays an important role in explaining the deviations between finite-time experiments and the theory for $N\rightarrow\infty$. It is present in the data of the mean return time $\langle n\rangle$ Fig. 3 and Fig. 4 as well as the dark states Fig. 5. By finite resolution, we mean that measurements are performed $N$ times per run. Typically, we record the target state for the first time in some random measurement attempt numbered $n_{i}\leq N$. We repeat this process by constructing an ensemble of trajectories to sample the mean return time which is maximally the number of runs. On the quantum computer that we choose the number of runs equal to $32,000$. We have already explained that in some cases we might not find the target within these $N$ measurements. Hence, in the figures presented, we show sample averages that, in principle, depend on our choice of $N$. More specifically, we define $K\leq 32,000$ as the number of paths for which we record the target state. Then, the sample mean $n$ conditioned on the target measurement state is $\langle n\rangle=\sum_{i=1}^{K}n_{i}\big{/}K$ where we exclude the cases that we did not detect. Far from the topological transitions, the option of zero detection per run (null measurement) of size $N$, is rare, and then the conditional measurement and the theoretical predictions yield similar results. However, as mentioned, close to topological transitions, we see deviations, which are due to the finite value of $N$. It should be noted that these deviations are a blessing in the sense that a precise measurement of a jump of $\langle n\rangle$, for a special value of $\tau$, as presented as black solid lines in Figs. 3 and 4, is non-physical, as the width of the transition cannot be zero. In the asymptotic theory ($N\rightarrow\infty$) for a special value of model parameters (here $\gamma\tau$ and $\alpha$), $\langle n\rangle$ exhibits a pointwise discontinuity related to a jump in the winding number and the dark lines in Fig. 5 become infinitely thin.
@@ VI.1 Broadening effect
We present an exact solution to the problem related to our experimental setup for both on-site and tracking measurement protocols. Previously, the broadening effect was studied in more generality for the on-site measurements, although the asymptotic behavior of large $N$ was studied there [39]. In the following we will discuss the broadening of the return time $\langle n\rangle$ for zero magnetic flux in detail.
VI.1.1 On-site protocol (broadening)
We first consider the broadening effect of the resonances of $\langle n\rangle$ using the on-site measurement protocol and $\alpha=0$. Let $F_{n}$ be the probability of finding the target for the first time at the $n-$th measurement. The case under study is exactly solvable [11, 40]. Specifically, with the eigenvalues of Eq. (3) we obtain
@@ VI.1.1 On-site protocol (broadening)
where $|z|^{2}=5/9+4/9\cos(3\gamma\tau)$. Here, we stick to the notation in [11], where in general, $z$ stands for the zeros of the generating function. When $|z|^{2}=1$, we have $F_{1}=1$ and then $\langle n\rangle=1$. Otherwise, $\langle n\rangle=\sum_{n=1}^{\infty}nF_{n}=2$, which is a behavior that was presented in Fig. 3 (c). Clearly, when $|z|^{2}$ is slightly less than one, so is $F_{1}$, yet from the recurrence theorem [5, 6] we have $\lim_{N\rightarrow\infty}\langle n(N)\rangle=2$ implying a very slow decay of $F_{n}\propto|z|^{2(n-2)}$ for $n>1$. Physically, close to $|z|^{2}=1$ we usually detect the system with the first click, but in rare events, we click no. Then the wave function does not overlap with the detected state, and recording a yes becomes unlikely. On average we find that the return click yes after $\langle n\rangle=2$ attempts which is remarkable since $F_{1}$ is nearly one.
@@
We will study this expression in the vicinity of the revival times, namely when $|z|^{2}\to 1$, where we also include the Zeno limit in this category. Note that if $|z|^{2}=1$ we have $\langle n(N)\rangle=1$ for any $N$.
We note that the expressions in Eqs. (11) and (12) have the same form for a two-level system, except for a different function of $|z|^{2}$ [40]. This indicates that for $\alpha=0$ the three-site model represents an effective two-level system.
@@
Recall that close to the transition point $\langle n\rangle=1$ for large $N$. Here by transition we mean a transition for $\langle n\rangle=2$ to $\langle n\rangle=1$ as we vary gamma tau across a resonance.
On the other hand, if we choose a control parameter like $\gamma\tau$ to be far from the transition, then $\langle n\rangle\approx 2$, since $N$ is large.
VI.1.2 Tracking protocol (broadening)
Now let us consider the broadening effect of the resonances of $\langle n\rangle$ using the tracking protocol and $\alpha=0$. The probability of finding the quantum walker the first time is